Exhibit 1 Total utility for hamburgers, fries, and Cokes Total Uitility from Hamburgers Total Utility from Fries Total Utility from Coke 1 hamburger (100 utils) 2 hamburgers (180 utils) 3 hamburgers (240 utils) 1 order of fries (30 utils) 1 Coke (40 utils) 2 orders of fries (50 utils) 2 Cokes (60 utils) 3 orders of fries (60 utils) 3 Cokes (70 utils) 3- In Exhibit 1 assume that the price of utility of having a second order of fries? hamburgers is $2 each, fries cost 50 cents each, and Cokes cost SI each. What is the marginal a. Ь. c, d. 10 utils. 20 utils. 30 utils. 50 utils. 4-In Exhibit 1, assume that the price of hamburgers is S2 each, fries cost 50 cents each, and Cokes cost SI each. Suppose the consumer has S6 to spend on hamburgers, fries, and Cokes. Which of the following meals gives the consumer the mast utility? a. 3 hamburgers, no fries, and no Cokes. b. 2 hamburgers, no fries, and 2 Cokes. c. 2 hamburgers, 2 orders of fries and 1 Coke. d. l hamburger, 2 orders of fries, and 3 Cokes. assume that the price of hamburgers is $2 each, fries cost 50 cents each, and Cokes cost S1 each. Suppose the mer has $6 each of the three goods? to spend on hamburgers, fries, and Cokes. In the consumer equilibrium, what is the marginal utility per dollar for
Exhibit 2 Production possibilities curve data Capital goods Consumption goods 25 23 19 19 13 0 6-Suppose an economyunits of consumption goods. economy is faced with the production possibilities t able shown in Exhibit 2. The first unit of capital goods will cost the a. b. 25 2 d. 23 7-Suppose an economy is faced with the production possibilities table shown in Exhibit 2. The second unit of capital goods production will cost-units of consumption goods, and the third unit ofcapital goods production will cost consumption goods units of 一一 4; 6 a. b. 25;23 c. 23; 19 d. e. 2;19 1; 23 8-Suppose an economy is faced with the production possibilities table shown in Exhibit 2 As additio produced, the opportunity cost in terms of sacrificed units of consumption goods because of of capital a. decreases; greater efficiency in production b. c. d. e. decreases; the law of increasing costs increases; decreasing opportunity cost increases; the law of increasing costs increases; greater efficiency in production
Answer #1

Solution:

EXHIBIT 1

Price of hamburgers, Ph = $2, price of fries, Pf = $0.50, price of coke, Pc = $1

Q3) Marginal utility is the additional satisfaction a person derives by consuming an extra unit. So, marginal utility of having a second order of fries is the additional utility the consumer receives by increasing the consumption of fries from one to two. So, in this case, marginal utility of second order of fries = TU2 - TU1 = 50 - 30 = 20 utils. (where TUi is total utility from order number i)

So, correct option is b) 20 utils.

Q4) Total income with the consumer, I = $6

So, the budget line of this consumer becomes: Ph*H + Pf*F + Pc*C = I (H, F, and C is the quantity of hamburgers, fries and coke respectively).

2*H + 0.5*F + 1*C = 6

For all the options, budget line is clearly satisfied. Going through each option individually,

Total utility from option (a): 240 + 0 + 0 = 240 utils

Total utility from option (b): 180 + 0 + 60 = 240 utils

Total utility from option (c): 180 + 50 + 40 = 270 utils

Total utility from option (d): 100 + 50 + 70 = 220 utils

Among all values, clearly 270 utils is highest of all. Thus, the meal giving most utility to the consumer is option (c) 2 H, 2 F and 1 C

Q5) From Q4, we have already established that where the consumer equilibrium is attained: 2 hamburgers, 2 orders of fries, and 1 coke. The marginal utility per dollar for a good i is the ratio of marginal utility attained at the equilibrium quantity of good i divided by the price of that good, that is,

Marginal utility per dollar = MUi/Pi

Marginal utility as explained in Q3, can be obtained at the equilibrium as follows:

Marginal utility of 2nd hamburger = 180 - 100 = 80 utils

Marginal utility of 2nd order of fries = 50 - 30 = 20 utils

Marginal utility of 1st coke = total utility of 1st coke (assuming that utility from consuming 0 coke is 0) = 40 utils

So, at equilibrium, marginal utility per dollar for hamburgers = 80/Ph = 80/2 = 40 utils

Marginal utility per dollar for fries = 20/Pf = 20/0.5 = 40 utils

Marginal utility per dollar for coke = 40/Pc = 40/1 = 40 utils

EXHIBIT 2:

Q6) As the economy moves from 0 unit of capital good to 1 unit of capital good, it automatically moves from 25 units of consumption goods to 23 units of consumption goods along the production possibility curve (PPC). So, the cost of first unit of capital good (that is the units of consumption goods economy will have to give up to get that first unit of capital good) = 25 - 23 = 2 units

Correct option: b) 2

Q7) Just as we solved above in Q6, we can find that the cost of:

Second unit of capital good = 23 - 19 = 4 units of consumption goods

Third unit of capital good = 19 - 13 = 6 units of consumption goods

So, the correct option is a) 4; 6

Q8) Clearly, from the above two questions we see that as the number of capital goods is increased in the economy, the quantity of consumption goods sacrificed increases (for 1st unit of capital good, cost is 2 units of consumption good. For second unit, it is 4 units of consumption goods; for third unit, cost is 6 units of consumption goods. So, clearly the sacrifice is increasing from 2 to 4 to 6 units) (eliminate options (a) and (e)). This cost of sacrifice is known as the opportunity cost, thus opportunity cost is increasing along the PPC. This happens due to decreased efficiency of each input towards the capital good, when moved away from the production of consumption good.

Thus, correct option is (c) increases; law of increasing costs